A block of mass 2 kg moving at a speed of 10 m/s accelerates at 3 m/s(2) for 5 sec. Compute its final kinetic energy.

Grade:11

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
here we make use of the formula
v=u+at
here u= 10 m/s
a= 3 m/s (square)
t= 5 sec
So, v= 25 m/s
therefor,
KE=\frac{1}{2}mv^{2}=\frac{1}{2}\times 2\times 25^{2}=625 Joules